OpenStudy (anonymous):

Evaluate the following. Write your answer in scientific notation. (3 x 107)(6.1 x 103) A. 18.3 x 1010 B. 18.3 x 1021 C. 1.83 x 1010 D. 1.83 x 1011

OpenStudy (anonymous):

why you can't use wolfram youself?

OpenStudy (anonymous):

A. 18.3 x 1010 = 18483 = 1.8483 x 10^4

OpenStudy (anonymous):

cause it doesnt give the answer :/

OpenStudy (anonymous):

thx Harkirat

OpenStudy (anonymous):

look it does, you just need to write ^ for power

OpenStudy (anonymous):

harkirat answer is wrong

OpenStudy (anonymous):

Why Tomas????

OpenStudy (anonymous):

where it says reslt its none of my answer choices

OpenStudy (anonymous):

*result

OpenStudy (anonymous):

how much r u writing?

OpenStudy (anonymous):

B. 18.3 x 1021 = 18684.3 = 1.86843 x 10^4

OpenStudy (anonymous):

because in scientific notation \[a⋅10^n\] a should be \[0≤a<10\]

OpenStudy (anonymous):

it's not B either

OpenStudy (anonymous):

Tomas pls check on what scientific notation is........

OpenStudy (anonymous):

check yourself http://en.wikipedia.org/wiki/Scientific_notation

OpenStudy (anonymous):

Lisa, are A, B, C, D questions themselves or are they options for answers?????

OpenStudy (anonymous):

In normalized scientific notation, the exponent b is chosen such that the absolute value of a remains at least one but less than ten (1 ≤ |a| < 10). Following these rules, 350 is always written as 3.5×102.

OpenStudy (anonymous):

also wolframalpha gives same answer, so wolframalpha and wikipedia both are incorrect?

OpenStudy (anonymous):

there the answer choices one of them is right

OpenStudy (anonymous):

tomas how did u get the right one the second time?

OpenStudy (anonymous):

\[3*6.1=18.3\] \[10^7*10^3=10^{10}\] \[18.3*10^{10}=1.83*10^{11}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx